OpenStudy (anonymous):

What are the vertices of the ellipse given by the equation x2 + 4y2 + 10x - 56y + 205 = 0?

OpenStudy (anonymous):

@Qawi Can you start with complete square ?

OpenStudy (anonymous):

i dont know how

OpenStudy (anonymous):

x² + 4y² + 10x - 56y + 205 = 0 ( x + 5) ² + 4 ( y - 7)² = 16

OpenStudy (anonymous):

Can you continue?